Could someone explain how I should solve these three problems;
Use the synthetic substitution to find the value of f(-2) for f(x)=x^3-3x^2-x-9.
Use the Factor Theorem to determine whether (x-3) is a factor of f(x)=x^4+12x^3+6x+27.
In x^3-4x^2+2x+1=0, how many possible positive and how many possible negative roots are there?

Answer by stanbon(75887)

Use the synthetic substitution to find the value of f(-2) for f(x)=x^3-3x^2-x-9.
---
-2)....1....-3....-1....-9
.........1....-5....9...|..-27
f(-2) = -27
===================================
Use the Factor Theorem to determine whether (x-3) is a factor of f(x)=x^4+12x^3+6x+27.
3)....1....12....0....6....27
........1....15...45...141..|..450
Since f(3) is not zero, x-3 is not a factor.
===================================
In x^3-4x^2+2x+1=0, how many possible positive and how many possible negative roots are there?
In f(x) there are 2 changes of sign, so 0 or 2 positive roots.
--------------
f(-x) = -x^3 - 4x^2 - 2x + 1
In f(-x) there are 2 changes of sign, so there are 0 or 2 negative roots.
===================================
Cheers,
Stan H.